### v1

# Quaternion-Domain Super MDS for Robust 3D Localization

Abstract
This paper proposes a novel low-complexity three-dimensional (3D) localization algorithm for wireless sensor networks, termed quanternion-domain super multi-dimensional scaling (QD-SMDS). The algorithm is based on a reformulation of the SMDS, originally developed in the real domain, using quaternion algebra. By representing 3D coordinates as quaternions, the method constructs a rank-$1$ Gram edge kernel (GEK) matrix that integrates both relative distance and angular information between nodes, which enhances the noise reduction effect achieved through low-rank truncation employing singular value decomposition (SVD), thereby improving robustness against information loss. To further reduce computational complexity, we also propose a variant of QD-SMDS that eliminates the need for the computationally expensive SVD by leveraging the inherent structure of the quaternion-domain GEK matrix. This alternative directly estimates node coordinates using only matrix multiplications within the quaternion domain. Simulation results demonstrate that the proposed method significantly improves localization accuracy compared to the original SMDS algorithm, especially in scenarios with substantial measurement errors. The proposed method also achieves comparable localization accuracy without requiring SVD.
I INTRODUCTION
With recent advancements in sensor technology, wireless sensor networks (WSNs) have increasingly emerged as fundamental information infrastructures across various industrial domains [1, 2, 3], including precision agriculture [4], smart factories [5], and medical sensing [6]. Among different types of sensor data, location information plays a particularly vital role—not only enhancing the value of the sensed data itself, but also contributing significantly to the efficient operation of WSNs. In fact, its importance is often regarded as comparable to that of payload data in conventional wireless communications. Many services utilizing such location information are designed under the assumption of large-scale networks consisting of numerous small, low-power sensor terminals (referred to as nodes[7, 8]. In these scenarios, there is a strong demand for algorithms capable of simultaneously estimating the positions of multiple nodes with high accuracy and low computational complexity, based on the aggregated multidimensional information available from WSNs [9, 10, 11].
Localization algorithms can be broadly categorized into three main approaches based on their underlying mathematical frameworks: Bayesian inference [12], convex optimization [13], and isometric embedding methods. When selecting an appropriate algorithm, it is crucial to consider the trade-off between computational complexity and estimation accuracy. This paper focuses on the isometric embedding approach, commonly known as multidimensional scaling (MDS) [14]. Unlike Bayesian inference and convex optimization, MDS offers the notable advantage of fixed computational complexity. While the convergence speed of Bayesian and optimization-based methods can vary significantly due to measurement noise or missing data, MDS maintains consistent complexity regardless of such factors. This makes MDS particularly well-suited for systems with limited computational resources and strict time constraints.
Among various MDS-based localization methods, super multi-dimensional scaling (SMDS) [16, 17] stands out as a hybrid algorithm that integrates both distance and angle information. It has demonstrated superior performance over classical MDS —which relies solely on distance information—even in the presence of angular uncertainties of approximately ±$\pm 35^{\circ}$  [18]. Complex-domain SMDS (CD-SMDS) is a technique that reduces computational complexity and improves precision by tailoring SMDS for two-dimensional (2D) localization [19]. Unlike conventional SMDS, which represents 2D node coordinates as real-valued vectors, CD-SMDS uses complex-valued scalars to express node positions. By constructing a GEK matrix in the complex domain that consolidates all measurement data, the method enables rank reduction to one and enhances accuracy through noise suppression via low-rank truncation using SVD [20]. This demonstrates that translating real-valued vectors into single complex scalars not only simplifies representation but also provides a foundation for extending SMDS to a 3D localization framework. In other words, if 3D node coordinates can similarly be expressed as scalars, it may be possible to develop an SMDS algorithm optimized for 3D localization.
As an intuitive candidate for representing 3D coordinates with a single scalar, one might initially consider bicomplex numbers [21]. However, to the best of our knowledge, no existing literature addresses SVD or matrix completion for bicomplex matrices, and their algebraic structure renders their integration into SMDS impractical. Accordingly, in this study, we turn our attention to quaternions as a mathematical tool for representing 3D coordinates.
Quaternions are an extension of complex numbers and are widely used to represent 3D rotations in fields such as computer graphics and computer vision  [27, 26]. A quaternion consists of one real component and three imaginary components  [22]. By utilizing three of its four degrees of freedom (DoF), it becomes possible to represent 3D node coordinates and construct a rank-$1$ GEK matrix in the quaternion domain. The algebraic structure of quaternions is well established, and frameworks for performing algebraic operations such as SVD are also available [23]. This enables us to exploit SVD-based low-rank truncation for noise suppression and thereby improve localization accuracy—leading to the development of QD-SMDS.
However, the computational efficiency gained by QD-SMDS is quite limited and may even result in increased computational cost in some systems. This is because performing SVD in the quaternion domain requires converting the target quaternion matrix into a complex-valued matrix of double the size, followed by applying SVD to the resulting complex-equivalent matrix. Given that the size of the GEK matrix increases proportionally with the number of node pairs, this process can become impractical—particularly in systems with constrained computational resources. To eliminate the need for SVD in SMDS, an iterative approach based on maximal ratio combining (MRC) was proposed in [19]. However, due to the non-commutative nature of quaternion multiplication, directly extending this method to QD-SMDS is not straightforward. In this paper, we address this challenge by explicitly considering quaternion non-commutativity in the combining operation and propose a novel, low-complexity 3D localization algorithm—termed quaternion-domain maximal ratio combining super multidimensional scaling (QD-MRC-SMDS) —which is constructed solely from simple quaternion multiplications. The contributions of this paper are summarized as follows [^1] The conference paper [24] is an earlier version of this paper, which was presented at the IEEE SPAWC 2025.:
  • A novel 3D localization algorithm, termed QD-SMDS is proposed. This algorithm is derived by reformulating the conventional SMDS, originally developed in the real domain, into the quaternion domain. By representing 3D coordinates using quaternions, a rank-1 GEK matrix can be constructed, enabling enhanced noise suppression through low-rank approximation. As a result, improved localization accuracy can be achieved, especially under large measurement errors.
  • To reduce the computational complexity, we propose a novel low-complexity variant, termed QD-MRC-SMDS. While QD-SMDS relies on SVD-based low-rank truncation, QD-MRC-SMDS performs localization using only simple multiplication operations derived from a closed-form expression. Furthermore, we extend this method into an iterative scheme that leverages the internal structure of the quaternion-domain GEK matrix, significantly improving estimation accuracy with only a marginal increase in computational cost.
  • To validate the efficacy of the proposed methods, we conducted computer simulations. As QD-SMDS requires additional measurement parameters compared to conventional SMDS, its performance was evaluated under various scenarios, including cases where these parameters are available, unavailable, or partially missing. The results demonstrate that QD-SMDS outperforms conventional SMDS, particularly under large angle measurement errors. Moreover, the computationally efficient variant achieves performance asymptotically close to that of QD-SMDS across nearly all scenarios.
[…]
II Preliminaries
In this section, we first provide a brief overview of quaternion algebra, and then introduce the quaternion singular value decomposition (QSVD).
II-A Basics of Quaternion Algebra
The quaternion space was first introduced by W. Hamilton  [25] as a natural extension of the complex space. A quaternion consists of one real component and three imaginary components, as follows:
$\mathbb{H}\triangleq\left\{a+\mathbf{i}b+\mathbf{j}c+\mathbf{k}d:a,b,c,d\right\},$ (1)
where $a,b,c,d\in\mathbb{R}$ are real numbers, and $\mathbf{i,j,k}$ are the imaginary units, which obey the following rules:
$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-1,$ (2)
with
[…]
If the real part of a quaternion is zero (i.e., $a=0$), the quaternion is called a pure quaternion. The square of a pure quaternion is the negative sum of the squares of its imaginary components:
$(\mathbf{i}b+\mathbf{j}c+\mathbf{k}d)^{2}=-b^{2}-c^{2}-d^{2},$ (4)
while multiplications in $\mathbb{H}$ is generally noncommutative, multiplication in $\mathbb{H}$ by real numbers is commutative, thus,
$\begin{array}[]{l}aq\neq qa\quad\mathrm{if}\;a\in\mathbb{H},\quad q\in\mathbb{H},\\ aq=qa\quad\mathrm{if}\;a\in\mathbb{R},\quad q\in\mathbb{H}.\end{array}$ (5)
For a quaternion $q=a+\mathbf{i}b+\mathbf{j}c+\mathbf{k}d$, its conjugate is defined as
$q^{*}=a-\mathbf{i}b-\mathbf{j}c-\mathbf{k}d.$ (6)
Conjugation in $\mathbb{H}$ shares similar properties with conjugation in $\mathbb{C}$. In $\mathbb{H}$, we have the following identities (assuming $q_{1}$ and $q_{2}$ are two arbitrary quaternions)
$\left(q_{1}+q_{2}\right)^{*}$ $=$ $q_{1}^{*}+q_{2}^{*},$ (7a)
$\left(q_{1}q_{2}\right)^{*}$ $=$ $q_{2}^{*}q_{1}^{*},$ (7b)
$\left(q_{1}^{*}\right)^{*}$ $=$ $q_{1}.$ (7c)
The norm of a quaternion $q$ is defined as
$\|q\|=\sqrt{qq^{*}}=\sqrt{q^{*}q}=\sqrt{a^{2}+b^{2}+c^{2}+d^{2}},$ (8)
while the reciprocal of a quaternion $q$ is defined as
$q^{-1}=\frac{q^{*}}{\|q\|^{2}}.$ (9)
II-B Quaternion Singular Value Decomposition (QSVD)
In this subsection, we introduce the QSVD as an algebraic tool to be used in the subsequent discussion. A quaternion matrix $\dot{\bm{Q}}\in\mathbb{H}^{M\times N}$ is written as $\dot{\bm{Q}}=\bm{Q}_{0}+\mathbf{i}\bm{Q}_{1}+\mathbf{j}\bm{Q}_{2}+\mathbf{k}\bm{Q}_{3}$, where $\bm{Q}_{l}\in\mathbb{R}^{M\times N}$ for $l=0,1,2,3$. Using the Cayley-Dickson notation [23], $\dot{\bm{Q}}$ can be expressed as
$\dot{\bm{Q}}=\bm{Q}_{\mathrm{a}}+\mathbf{j}\bm{Q}_{\mathrm{b}},$ (10)
where $\bm{Q}_{\mathrm{a}}=\bm{Q}_{0}+\mathbf{i}\bm{Q}_{1}\in\mathbb{C}^{M\times N},\bm{Q}_{\mathrm{b}}=\bm{Q}_{2}+\mathbf{i}\bm{Q}_{3}\in\mathbb{C}^{M\times N}$.
Accordingly, the equivalent complex-valued representation of the quaternion matrix $\dot{\bm{Q}}$ is given by
$\bm{Q}_{\mathrm{c}}=\begin{bmatrix}\bm{Q}_{\mathrm{a}}&\bm{Q}_{\mathrm{b}}\\ -\bm{Q}_{\mathrm{b}}^{*}&\bm{Q}_{\mathrm{a}}^{*}\end{bmatrix}\in\mathbb{C}^{2M\times 2N}.$ (11)
The SVD of the quaternion matrix $\dot{\bm{Q}}$ can be derived from the SVD of its equivalent complex matrix $\bm{Q}_{\mathrm{c}}$. Given $\bm{\dot{Q}}=\bm{U}_{\mathrm{q}}\bm{D}_{\mathrm{q}}\bm{V}_{\mathrm{q}}^{\mathsf{H}}$ and $\dot{Q}_{\mathrm{c}}=\bm{U}_{\mathrm{c}}\bm{D}_{\mathrm{c}}\bm{V}_{\mathrm{c}}^{\mathsf{H}}$, we have
$\bm{D}_{\mathrm{q}}$ $=$ $\mathrm{r}_{\mathrm{odd}}\left[\mathrm{c}_{\mathrm{odd}}\left[\bm{D}_{\mathrm{c}}\right]\right],$ (12a)
$\bm{U}_{\mathrm{q}}$ $=$ $\mathrm{c}_{\mathrm{odd}}\left[\bm{U}_{1}\right]+\mathbf{j}\mathrm{c}_{\mathrm{odd}}\left[-\bm{U}_{2}^{*}\right],$ (12b)
$\bm{V}_{\mathrm{q}}$ $=$ $\mathrm{c}_{\mathrm{odd}}\left[\bm{V}_{1}\right]+\mathbf{j}\mathrm{c}_{\mathrm{odd}}\left[-\bm{V}_{2}^{*}\right],$ (12c)
where
$\bm{U}_{\mathrm{c}}=\begin{bmatrix}\bm{U}_{1}\\ \bm{U}_{2}\end{bmatrix},\quad\bm{V}_{\mathrm{c}}=\begin{bmatrix}\bm{V}_{1}\\ \bm{V}_{2}\end{bmatrix},$ (13)
with $\bm{U}_{1},\bm{U}_{2}\in\mathbb{C}^{M\times 2M}$ and $\bm{V}_{1},\bm{V}_{2}\in\mathbb{C}^{N\times 2N}$.
III Problem Formulation and the SMDS Algorithm
III-A Problem Formulation
Consider a network embedded in 3D Euclidean space consisting of $N$ nodes, among which $N_{\mathrm{A}}$ nodes are designated as anchor nodes (ANs) with known, error-free locations. The remaining $N_{\mathrm{T}}\triangleq N-N_{\mathrm{A}}$ nodes, referred to as target nodes (TNs), have unknown locations to be estimated. It is assumed that relative distances and angles are measurable between any pair of ANs and between each ANTN pair, while such measurements are unavailable among TNs. The objective of this paper is to estimate the coordinates of TNs based on the measured (and noisy) relative distances and angles between nodes, as well as the known coordinates of ANs.
Let the coordinates of the $n$-th node in the network be denoted by the column vector $\bm{x}\triangleq\left[a_{n},b_{n},c_{n}\right]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$, which represents the 3D coordinates of the node in the Cartesian coordinate system. We define the coordinate matrix $\bm{X}_{\mathrm{A}}$ as the matrix that stacks the coordinate vectors of ANs:
$\bm{X}_{\mathrm{A}}\triangleq\left[\bm{x}_{1},\ldots,\bm{x}_{N_{\mathrm{A}}}\right]^{\mathsf{T}}\in\mathbb{R}^{N_{\mathrm{A}}\times 3},$ (14)
and similarly define the coordinate matrix that stacks the coordinate vectors of TNs as
$\bm{X}_{\mathrm{T}}\triangleq\left[\bm{x}_{1},\ldots,\bm{x}_{N_{\mathrm{T}}}\right]^{\mathsf{T}}\in\mathbb{R}^{N_{\mathrm{T}}\times 3}.$ (15)
The real-valued matrix carrying the coordinate vectors of all nodes in the network can then be expressed as
$\bm{X}\triangleq\left[\bm{X}_{\mathrm{A}}^{\mathsf{T}},\bm{X}_{\mathrm{T}}^{\mathsf{T}}\right]^{\mathsf{T}}\in\mathbb{R}^{N\times 3}.$ (16)
Consider the set $\mathcal{M}$ of unique index pairs $(i,j)$, arranged in ascending order, for which the pairwise distances and phases are measurable, i.e., any pair among ANs or between ANs and TNs:
$\mathcal{M}\triangleq\left\{(1,2),\cdots,(1,N),(2,3),\cdots,(2,N),\cdots,(N_{\mathrm{A}},N)\right\},$ (17)
such that each pair $m\in\mathcal{M}$ corresponds to an edge vector $\bm{v}_{m}$ is defined as
$\bm{v}_{m}=\bm{x}_{i}-\bm{x}_{j},\qquad j>i.$ (18)
From the above, the real-valued edge matrix consisting of the collection of all $M\triangleq|\mathcal{M}|=N_{\mathrm{A}}(N_{\mathrm{A}}-1)/2+N_{\mathrm{A}}N_{\mathrm{T}}$ edges is defined by
$\bm{V}\triangleq\left[\bm{v}_{1},\ldots,\bm{v}_{m},\ldots,\bm{v}_{M}\right]^{\mathsf{T}}=\bm{C}\bm{X}\in\mathbb{R}^{M\times 3},$ (19)
where $\bm{C}\triangleq\left[\bm{C}_{\mathrm{AA}},\bm{C}_{\mathrm{AT}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times N}$ is a structure matrix encoding the pairwise relationships between nodes and edges.
The submatrix corresponding to the edges between ANs, denoted as $\bm{C}_{\mathrm{AA}}\in\mathbb{R}^{N_{\mathrm{A}}(N_{\mathrm{A}}-1)/2\times N}$, is expressed as
[…]
while the submatrix corresponding to the edges between ANs and TNs, denoted as $\bm{C}_{\mathrm{AT}}\in\mathbb{R}^{N_{\mathrm{A}}N_{\mathrm{T}}\times N}$, is expressed as
[…]
III-B SMDS Algorithm Recap
In this subsection, we briefly describe the conventional SMDS algorithm [16, 17] for estimating the coordinates of TNs based on the above formulation.
The inner product between two edge vectors $\bm{v}_{m}$ and $\bm{v}_{p}$ ($m,p\in\mathcal{M}$) can be expressed as
$k_{mp}^{\mathrm{r}}\triangleq\left<\bm{v}_{m},\bm{v}_{p}\right>=d_{m}d_{p}\cos\alpha_{mp},$ (21)
where $d_{m}\triangleq\|\bm{v}_{m}\|$ denotes the Euclidean distance between the two nodes (i.e., $\bm{x}_{i}$ and $\bm{x}_{j}$), and $\alpha_{mp}$ denotes the angle difference of arrival (ADoA) between $\bm{v}_{m}$ and $\bm{v}_{p}$.
Based on (21), the real-domain GEK matrix $\bm{K}_{\mathrm{r}}\in\mathbb{R}^{M\times M}$, which integrates both relative distance and ADoA information, can be expressed as
$\bm{K}_{\mathrm{r}}\!\!\!$$\triangleq$ $\!\!\!\bm{V}\bm{V}^{\mathsf{T}}$ (22)
$=$ $\!\!\!\mathrm{diag}\left(\bm{d}\right)\begin{bmatrix}\cos\alpha_{11}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{1M}\\ \vdots&\ddots&\vdots\\ \cos\alpha_{M1}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left(\bm{d}\right),$
where $\bm{d}=\left[d_{1},\ldots,d_{m},\ldots,d_{M}\right]^{\mathrm{T}}\in\mathbb{R}^{M\times 1}$.
Algorithm 1 SMDS Algorithm
1:Input:
2:Measured pairwise distances and ADoAs: $\tilde{d}_{m}$ and $\tilde{\alpha}_{mp}$.
3:The coordinates of at least $4$ ANs.
4:Steps:
5:Construct the real-domain GEK matrix $\tilde{\bm{K}}_{\mathrm{r}}$ in Eq. (22) using the input parameters.
6:Perform SVD of the constructed GEK matrix $\tilde{\bm{K}}_{\mathrm{r}}$.
[…]
8:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ in Eq. (24).
9:Apply the Procrustes transform to $\hat{\bm{X}}$ if needed (e.g., see [28]).
As shown in (22), it is evident that the rank of the real-domain GEK matrix is $3$.
Assuming that all pairwise distance and ADoA measurements required in (22) are available, the real-domain GEK matrix with measurement errors, denoted as $\tilde{\bm{K}}_{\mathrm{r}}$, can be constructed. Given the SVD of the real-valued GEK matrix $\tilde{\bm{K}}_{\mathrm{r}}$ as $\tilde{\bm{K}}_{\mathrm{r}}=\bm{U}\bm{\varLambda}\bm{U}^{\mathsf{T}}$, the estimate of the edge matrix $\hat{\bm{V}}$ is then obtained as
$\hat{\bm{V}}=\bm{U}_{M\times 3}\bm{\Lambda}_{3\times 3}^{\odot\frac{1}{2}},$ (23)
where $\bm{U}_{\bm{M}\times 3}$ consists of the first $3$ columns of $\bm{U}$, and $\bm{\Lambda}_{3\times 3}$ contains the corresponding top $3$ singular values on its diagonal.
Finally, the estimated coordinate matrix can be recovered from $\hat{\bm{V}}$ by inverting the relationship in (19), i.e., $\hat{\bm{X}}=\bm{C}^{-1}\hat{\bm{V}}$. However, since the rank of the structure matrix $\bm{C}$, as defined in (20), is $N-1$, this inverse cannot be directly computed. To circumvent this rank deficiency, we exploit the known coordinate matrix $\bm{X}_{\mathrm{A}}$ corresponding to ANs in (14). By incorporating this prior knowledge, the inversion can be reformulated as
$\left[\begin{array}[]{c}\bm{X}_{\mathrm{A}}\\ \hline\cr\hat{\bm{X}}\end{array}\right]=\left[\begin{array}[]{c|c}\bm{I}_{N_{\mathrm{A}}}&\bm{0}_{N_{\mathrm{A}}\times N_{\mathrm{T}}}\\ \hline\cr\bm{0}_{M\times N_{A}}&\bm{C}\end{array}\right]^{-1}\left[\begin{array}[]{c}\bm{X}_{\mathrm{A}}\\ \hline\cr\hat{\bm{V}}\end{array}\right].$ (24)
Moreover, since the SMDS algorithm operates solely on the relative relationships among nodes, the inverse problem in (24) can be characterized in multiple ways, resulting in a non-unique solution. Therefore, a Procrustes transformation [28] is typically required to align the estimated $\hat{\bm{X}}$ with the true coordinates $\bm{X}$ in terms of scale, orientation, and translation.
For the sake of completeness, this subsection is concluded with the pseudo-code of the full SMDS algorithm, presented in Algorithm 1.
IV Quaternion-Domain Super MDS
(a) Cartesian coordinate system
(b) $(\mathrm{x},\mathrm{y})$-plane
(c) $(\mathrm{x},\mathrm{z})$-plane
(d) $(\mathrm{y},\mathrm{z})$-plane
Figure 1: Illustration of the parameters required to construct quaternion-domain GEK matrix $\bm{K}_{\mathrm{q}}$.
IV-A QD-SMDS Algorithm
IV-A1 Derivation of the Proposed Algorithm
In this subsection, we derive QD-SMDS —an extension of the conventional SMDS tailored for 3D localization—using the quaternion algebra introduced in Section II.
First, the 3D coordinates vector $\bm{x}_{n}$ of a generic node $n$ in the network can be alternatively expressed by the quaternion representation $\chi\in\mathbb{H}$ as
$\bm{x}_{n}=\left[a_{n},b_{n},c_{n}\right]\Longleftrightarrow\chi_{n}=a_{n}+\mathbf{i}b_{n}+\mathbf{j}c_{n}+\mathbf{k}\cdot 0,$ (25)
where out of the four DoF, three are used for the $(x,y,z)$ coordinates, and the remaining one is set to 00222In this paper, the component corresponding to the imaginary unit $\mathbf{k}$ is set to zero. Although any one of the four components can be chosen as zero, it should be noted that the GEK matrix must be modified accordingly if a different component is selected. .
Accordingly, the quaternion coordinate vector corresponding to the real coordinate matrix defined in (16) can be expressed as
$\bm{\chi}\triangleq\left[\chi_{1},\ldots,\chi_{N}\right]\in\mathbb{H}^{N\times 1}.$ (26)
Similarly, the edge vector $\bm{v}_{m}$ between any two nodes $\bm{x}_{i}$ and $\bm{x}_{j}$ in (18) can be represented as
$\nu_{m}=(\underbrace{a_{i}-a_{j}}_{\grave{a}_{m}})+\mathbf{i}(\underbrace{b_{i}-b_{j}}_{\grave{b}_{m}})+\mathbf{j}(\underbrace{c_{i}-c_{j}}_{\grave{c}_{m}})+\mathbf{k}\cdot 0,$ (27)
where $\grave{a}_{m}\triangleq a_{i}-a_{j},\grave{b}_{m}\triangleq b_{i}-b_{j},\grave{c}_{m}\triangleq c_{i}-c_{j}$. From the above, the quaternion edge vector corresponding to the real edge matrix defined in (19) can be written as
$\bm{\nu}\triangleq\left[\nu_{1},\ldots,\nu_{m},\ldots,\nu_{M}\right]^{\mathsf{T}}=\bm{C}\bm{\chi}\in\mathbb{H}^{M\times 1},$ (28)
with the structure matrix $\bm{C}$ defined in (20).
Next, the QD-SMDS algorithm is derived based on the quaternion formulation given in (25)–(28). In this context, the inner product defined in (21) can be rewritten as
$\bm{v}_{m},\bm{v}_{p}\right>\triangleq\left[\grave{a}_{m},\grave{b}_{m},\grave{c}_{m}\right]\begin{bmatrix}\grave{a}_{p}\\ \grave{b}_{p}\\ \grave{c}_{p}\end{bmatrix}$
$\!\!\!\!\quad=\grave{a}_{m}\grave{a}_{p}+\grave{b}_{m}\grave{b}_{p}+\grave{c}_{m}\grave{c}_{p}=d_{m}d_{p}\cos\alpha_{mp}.$ (29)
In turn, the outer product of the vectors obtained by projecting the edge vectors onto the $(\mathrm{x},\mathrm{y})$ plane can be expressed as
$\!\!\!\!\left|\bm{v}_{m}^{(\mathrm{xy})}\times\bm{v}_{p}^{(\mathrm{xy})}\right|\triangleq\mathrm{det}\left[\begin{bmatrix}\grave{a}_{m}\quad\grave{b}_{m}\\ \grave{a}_{p}\quad\grave{b}_{p}\end{bmatrix}\right]$
$\!\!\!\!\quad=\grave{a}_{m}\grave{b}_{p}-\grave{a}_{p}\grave{b}_{m}=d_{m}^{(\mathrm{xy})}d_{p}^{(\mathrm{xy})}\sin\alpha_{mp}^{(\mathrm{xy})}.$ (30a)
Similarly, the outer products on the $(\mathrm{x},\mathrm{z})$ and $(\mathrm{y},\mathrm{z})$ planes, respectively, can be expressed as
$\left|\bm{v}_{m}^{(\mathrm{xz})}\times\bm{v}_{p}^{(\mathrm{xz})}\right|=\grave{a}_{m}\grave{c}_{p}-\grave{a}_{p}\grave{c}_{m}=d_{m}^{(\mathrm{xz})}d_{p}^{(\mathrm{xz})}\sin\alpha_{mp}^{(\mathrm{xz})},$ (30b)
$\left|\bm{v}_{m}^{(\mathrm{yz})}\times\bm{v}_{p}^{(\mathrm{yz})}\right|=\grave{b}_{m}\grave{c}_{p}-\grave{b}_{p}\grave{c}_{m}=d_{m}^{(\mathrm{yz})}d_{p}^{(\mathrm{yz})}\sin\alpha_{mp}^{(\mathrm{yz})},$ (30c)
where $\bm{v}_{m}^{(\mathrm{xy})}$, $\bm{v}_{m}^{(\mathrm{xz})}$, and $\bm{v}_{m}^{(\mathrm{yz})}$ denote the 2D vectors obtained by projecting $\bm{v}_{m}$ onto the $(\mathrm{x},\mathrm{y})$, $(\mathrm{x},\mathrm{z})$, and $(\mathrm{y},\mathrm{z})$ planes, respectively.
The corresponding Euclidean norms (distances) are give by $d_{m}^{(\mathrm{xy})}\triangleq\|\bm{v}_{m}^{(\mathrm{xy})}\|$, $d_{m}^{(\mathrm{xz})}\triangleq\|\bm{v}_{m}^{(\mathrm{xz})}\|$, and $d_{m}^{(\mathrm{yz})}\triangleq\|\bm{v}_{m}^{(\mathrm{yz})}\|$. Similarly, $\alpha_{mp}^{(\mathrm{xy})}$, $\alpha_{mp}^{(\mathrm{xz})}$, and $\alpha_{mp}^{(\mathrm{yz})}$ represent the ADoAs between two 2D projected vectors on the $(\mathrm{x},\mathrm{y})$, $(\mathrm{x},\mathrm{z})$, and $(\mathrm{y},\mathrm{z})$ planes, respectively. For clarity of these parameters, please refer to Fig.1 at the top of this page.
Based on (29)–(30), the product of the quaternion edges $\nu_{m}$ and $\nu_{p}^{*}$, with $m\neq p$, can be expressed as
$\nu_{m}\nu_{p}^{*}$ $=\underbrace{\left(\grave{a}_{m}\grave{a}_{p}+\grave{b}_{m}\grave{b}_{p}+\grave{c}_{m}\grave{c}_{p}\right)}_{\left<\bm{v}_{m},\bm{v}_{p}\right>}+\mathbf{i}\underbrace{\left(\grave{a}_{p}\grave{b}_{m}-\grave{a}_{m}\grave{b}_{p}\right)}_{-\left|\bm{v}_{m}^{(\mathrm{xy})}\times\bm{v}_{p}^{(\mathrm{xy})}\right|}$ (31)
$\quad+\mathbf{j}\underbrace{\left(\grave{a}_{p}\grave{c}_{m}-\grave{a}_{m}\grave{c}_{p}\right)}_{-\left|\bm{v}_{m}^{(\mathrm{xz})}\times\bm{v}_{p}^{(\mathrm{xz})}\right|}+\mathbf{k}\underbrace{\left(\grave{b}_{p}\grave{c}_{m}-\grave{b}_{m}\grave{c}_{p}\right)}_{-\left|\bm{v}_{m}^{(\mathrm{yz})}\times\bm{v}_{p}^{(\mathrm{yz})}\right|}$
$=d_{m}d_{p}\cos\alpha_{mp}-\mathbf{i}d_{m}^{(\mathrm{xy})}d_{p}^{(\mathrm{xy})}\sin\alpha_{mp}^{(\mathrm{xy})}$
$\quad-\mathbf{j}d_{m}^{(\mathrm{xz})}d_{p}^{(\mathrm{xz})}\sin\alpha_{mp}^{(\mathrm{xz})}-\mathbf{k}d_{m}^{(\mathrm{yz})}d_{p}^{(\mathrm{yz})}\sin\alpha_{mp}^{(\mathrm{yz})}\!.$
$\bm{K}_{\mathrm{q}}\triangleq\bm{\nu}\bm{\nu}^{\mathsf{H}}=\mathrm{diag}\left[\bm{d}\right]\begin{bmatrix}\cos\alpha_{11}&\cdots&\cos\alpha_{1M}\\ \vdots&\ddots&\vdots\\ \cos\alpha_{M1}&\cdots&\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left[\bm{d}\right]-\mathbf{i}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xy})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xy})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]$
$\quad-\mathbf{j}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]-\mathbf{k}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{yz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{yz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right].$ (32)
From (31), the rank-$1$ quaternion-domain GEK matrix, which incorporates all pairwise distances and ADoA information among the nodes, is given in (IV-A1), shown at the top of the next page. Assuming that measured values of all pairwise distance and ADoA parameters appearing in (31) are available, the quaternion-domain GEK matrix with measurement errors, denoted as $\tilde{\bm{K}}_{\mathrm{q}}$, can be constructed. Accordingly, the estimate of the quaternion edge vector $\bm{\nu}$ is given as
$\hat{\bm{\nu}}=\sqrt{\lambda}\bm{u},$ (33)
where $(\lambda,\bm{u})$ denotes the pair consisting of the largest eigenvalue and its corresponding eigenvector of $\tilde{\bm{K}}_{\mathrm{q}}$.
The SVD of $\tilde{\bm{K}}_{\mathrm{q}}$ is performed using the QSVD introduced in Section II-B. Note that, in our case, since the quaternion-domain GEK $\tilde{\bm{K}}_{\mathrm{q}}$ is regular, the quaternion eigendecomposition is effectively equivalent to the QSVD.
In the conventional SMDS algorithm, the GEK matrix is constructed based on the real-valued vectors described in (18). As a result, its rank is limited to $3$, as clearly shown in (23), which inherently restricts the noise reduction capability achievable through low-rank truncation via SVD. In contrast, the proposed QD-SMDS algorithm utilizes a quaternion-domain representation, enabling the GEK matrix to have rank-$1$. This low-rank structure allows for maximal noise reduction, thereby significantly enhancing the robustness of localization performance against measurement errors.
Finally, we estimate the real-valued coordinate matrix $\bm{X}$ in (16) from the estimated quaternion edge vector $\hat{\bm{\nu}}$. First, the real, $\bm{\mathrm{i}}$-, and $\bm{\mathrm{j}}$-components of $\hat{\bm{\nu}}$ are extracted and rearranged according to the $(\mathrm{x},\mathrm{y},\mathrm{z})$ coordinates based on (27), yielding the estimated real-valued edge matrix as
$\hat{\bm{V}}\triangleq\left[\hat{\bm{v}}_{1},\cdots,\hat{\bm{v}}_{M}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times 3}.$ (34)
Given the estimated edge vector matrix $\hat{\bm{V}}$, the same procedure as in (24) can be applied to obtain the estimated coordinate matrix $\hat{\bm{X}}$. Finally, a Procrustes transformation is applied to $\hat{\bm{X}}$ using the known positions of ANs.
IV-A2 Construction of the Quaternion-Domain GEK Matrix
To construct the quaternion-domain GEK matrix in (IV-A1), additional phase difference information is required—beyond the typically measurable pairwise distance $d_{m}$ and ADoA $\alpha_{mp}$—when the positional relationship between nodes is projected onto each plane. Depending on the extent to which this additional information can be obtained from measurements, two practical scenarios are considered.
The first scenario, referred to as Scenario I, assumes that only the mutual distances $d_{m}$ and ADoAs $\alpha_{mp}$ between nodes are available, with no additional angular information. Under this condition, only the real-domain GEK matrix used in the conventional SMDS algorithm, as defined in (22), can be directly constructed from measurements; the quaternion-domain GEK cannot. Therefore, it is necessary to first execute the conventional SMDS to estimate the coordinates of TNs. Based on these estimated positions, the angular information required for constructing the quaternion-domain GEK matrix can then be computed. With this information in place, the QD-SMDS algorithm can subsequently be executed to refine the positioning accuracy.
The second scenario, referred to as Scenario II, assumes that azimuth and elevation angles can be measured by using planar antennas, as proposed in [29], deployed at each AN. By appropriately orienting the planar antennas, it becomes possible to obtain the parameters illustrated in Fig. 2, where $\theta$ denotes the elevation angle and $\phi$ denotes the azimuth angle. In the figure, red, blue, and green colors indicate the parameters obtained by projecting the edge vectors onto the $(\mathrm{x},\mathrm{y})$, $(\mathrm{x},\mathrm{z})$, and $(\mathrm{y},\mathrm{z})$ planes, respectively.
From the above, the quaternion-domain GEK matrix $\bm{K}_{\mathrm{q}}$ can be constructed based on all measurable all pairwise distances and angular information:
$d_{m},\theta_{m}^{(\mathrm{x})},\theta_{m}^{(\mathrm{y})},\theta_{m}^{(\mathrm{z})},\phi_{m}^{(\mathrm{xy})},\phi_{m}^{(\mathrm{yz})},\phi_{m}^{(\mathrm{xz})},\alpha_{mp},$ (35)
where the following parameters must be computed in advance from the quantities listed in (35) as
$d_{m}^{(\mathrm{xy})}$ $=$ $d_{m}\sin\theta_{m}^{(\mathrm{z})},$ (36a)
$d_{m}^{(\mathrm{xz})}$ $=$ $d_{m}\sin\theta_{m}^{(\mathrm{y})},$ (36b)
$d_{m}^{(\mathrm{yz})}$ $=$ $d_{m}\sin\theta_{m}^{(\mathrm{x})},$ (36c)
$\alpha_{mp}^{(\mathrm{xy})}$ $=$ $\phi_{p}^{(\mathrm{xy})}-\phi_{m}^{(\mathrm{xy})},$ (36d)
$\alpha_{mp}^{(\mathrm{xz})}$ $=$ $\phi_{p}^{(\mathrm{xz})}-\phi_{m}^{(\mathrm{xz})},$ (36e)
$\alpha_{mp}^{(\mathrm{yz})}$ $=$ $\phi_{p}^{(\mathrm{yz})}-\phi_{m}^{(\mathrm{yz})}.$ (36f)
Figure 2: Parameters that can be obtained using a planar antenna.
Finally, we conclude this subsection by summarizing the QD-SMDS algorithm in the form of pseudo-code, presented in Algorithm 2 at the top of the next page.
Algorithm 2 Quaternion-Domain SMDS
1:Input:
2:Measured and estimated pairwise distances and ADoAs: $d_{m},d_{m}^{(\mathrm{xy})},d_{m}^{(\mathrm{xz})},d_{m}^{(\mathrm{yz})},\alpha_{mp},\alpha_{mp}^{(\mathrm{xy})},\alpha_{mp}^{(\mathrm{xz})},\alpha_{mp}^{(\mathrm{yz})}$
3:The coordinates of at least 4 ANs.
4:Steps:
5:Construct the quaternion-domain GEK matrix $\tilde{\bm{K}}$ in (IV-A1) using the input parameters.
6:Perform QSVD of the constructed GEK matrix $\tilde{\bm{K}}$ (see Section II-II-B).
7:Obtain the edge vector $\hat{\bm{\nu}}$ using Eq. (33).
8:Convert the estimated quaternion edge vector $\hat{\bm{\nu}}$ to the estimated real-valued edge matrix $\hat{\bm{V}}$.
9:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ using Eq. (24).
10:Apply the Procrustes transform to $\hat{\bm{X}}$ if needed (e.g., see [28]).
IV-B Handling Incomplete Data
In the previous sections, it was assumed that all measurements between ANs and TNs were fully available, i.e., obtained without any missing data. However, in practical scenarios, partial information loss may occur due to environmental factors such as non-line-of-sight (NLOS) conditions. In such cases, it becomes necessary to apply a matrix completion algorithm prior to executing the SMDS or QD-SMDS algorithms.
A typical example is the scenario in which some of the pairwise distance measurements in the network are missing. In such case, the Euclidean distance matrix (EDM) completion algorithm proposed in  [30] can be employed to recover the missing entries. Consequently, most studies assume that the pairwise distances are either directly measured or completed in advance.
Another practical scenario arises when the angle measurements in the network are partially unavailable, resulting in a partially observed GEK matrix. In this case, two different completion approaches can be considered. The first approach leverages the fact that the EDM is fully observed. Classical MDS algorithm can be applied to estimate the coordinates of TNs, from which the missing angle information can be inferred. The second approach exploits the sparsity of the GEK matrix itself. A matrix completion method based on low-rank approximation can be directly applied to the sparse GEK matrix. For the real-domain GEK matrix used in the SMDS algorithm, the method proposed in [31] is sufficient. However, this algorithm cannot be directly applied to the quaternion-domain GEK matrix in QD-SMDS. To address this, advanced quaternion matrix completion algorithms, such as those proposed in [23], can be employed to recover the missing entries. As demonstrated in the simulation results presented later, the QD-SMDS algorithm tends to outperform the SMDS algorithm in scenarios where angle measurements are partially unavailable. This advantage arises from the fact that low-rank matrix completion methods generally yield better results when applied to large-scale matrices with lower rank. While both SMDS and QD-SMDS construct GEK of size $M\times M$, the rank of the matrix in SMDS is $3$, whereas in QD-SMDS it is only $1$. This lower rank facilitates more accurate completion, thereby enhancing localization performance. Even when using classical matrix completion methods based on the nuclear norm, such as the one proposed in [31], the QD-SMDS algorithm is still expected to outperform the SMDS algorithm. Although these methods cannot be directly applied to quaternion matrices, the quaternion-domain GEK matrix can be decomposed into two complex-valued matrices, $\bm{Q}_{\mathrm{a}}$ and $\bm{Q}_{\mathrm{b}}$, via the Cayley-Dickson construction in (10). Matrix completion is then performed separately on these two complex matrices. Since each has rank $2$, the completion tends to be more effective than applying the same algorithm directly to the rank-$3$ GEK matrix in the SMDS algorithm. In other words, QD-SMDS inherently exhibits greater robustness to missing data due to its structural properties, as shown in the following section.
IV-C Performance Assessment
IV-C1 Simulation Conditions
Computer simulations were conducted to evaluate the performance of the proposed QD-SMDS algorithm. The simulation environment assumes a room with a dimensions of $30$[m] (length) $\times$ $30$[m] (width) $\times$ $10$[m] (height). ANs were placed at five locations: the four upper corners of the room, specifically at $(x,y,z)=(0,0,10),(30,0,10),(30,30,10),$ and $(0,30,10)$, as well as the origin $(x,y,z)=(0,0,0)$. TNs were randomly placed at 15 locations within the interior, with their $\mathrm{x},\mathrm{y}$, and $\mathrm{z}$ coordinates independently drawn from a uniform distribution.
Distance measurements are modeled as Gamma-distributed random variables [32] with the mean equal to the true distance $d$ and a standard deviation of $\sigma_{d}$. The probability density function (PDF) of the measured distances $\tilde{d}$ corresponding to the true distance $d$ is given by
$p_{\mathrm{D}}(d;\alpha.\beta)=\left(\beta^{\alpha}\Gamma(\alpha)\right)^{-1}\tilde{d}^{\left(\alpha-1\right)}e^{\frac{\tilde{d}}{\beta}}.\,$ (37)
where $\alpha\triangleq d^{2}/\sigma^{2}_{d}$ and $\beta\triangleq\sigma^{2}_{d}/d$.
In turn, angle measurement errors $\delta_{\theta}$ are assumed to follow a Tikhonov-distribution [33, 34]. The PDF of the measured angle $\tilde{\theta}=\theta+\delta_{\theta}$ corresponding to the true angle $\theta$ is given by
$p_{\mathrm{\Theta}}\left(\tilde{\theta};\theta,\rho\right)=\frac{1}{2\pi I_{0}(\rho)}\exp\left[\rho\cos(\theta-\tilde{\theta})\right].$ (38)
The range of the angular error is determined by the angular parameter $\epsilon$, which represents the bounding angle of the central $90^{\mathrm{th}}$ percentile and is expressed as
[…]
Estimation errors are evaluated using the Frobenius norm of the difference between the estimated TN coordinate matrix $\hat{\bm{X}}_{\mathrm{T}}$ and the true coordinate matrix $\bm{X}_{\mathrm{T}}$, i.e.,
$\xi\triangleq\frac{1}{N_{t}}\|\hat{\bm{X}}_{\mathrm{T}}-\bm{X}_{\mathrm{T}}\|_{\mathrm{F}}.$ (40)
Figure 3: Comparison of localization accuracy between the SMDS and QD-SMDS algorithms in Scenario I.
Figure 4: Comparison of localization accuracy between the SMDS and QD-SMDS algorithms in Scenario II.
IV-C2 Simulation Results Without Missing Data
Based on the above simulation conditions, we have compared the localization accuracy of the conventional SMDS and the proposed QD-SMDS algorithms under both Scenario I and Scenario II described in Section IV-A. Fig. 3 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario I. The horizontal axis represents the standard deviation of the Gamma-distributed distance measurements, as defined in (37), while the vertical axis indicates the averaged estimation error, as defined in (40). The results are plotted for various angular measurement errors $\epsilon\in\left\{10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ}\right\}$.
When the angle error is small ($\epsilon=10^{\circ}$ and $20^{\circ}$), we can observe that the relative performance of the two methods varies depending on the distance error. The QD-SMDS algorithm outperforms the SMDS algorithm up to $\sigma_{d}=1.0$ [m] for $\epsilon=10^{\circ}$, and up to $\sigma_{d}=1.8$ [m] for $\epsilon=20^{\circ}$. However, as the distance error increases further, SMDS begins to achieve higher accuracy. This behavior can be attributed to the fact that when angle errors are minimal, the GEK matrix can be constructed with high precision, making aggressive noise suppression via low-rank truncation less critical. Furthermore, SMDS, which requires fewer (and noisy) parameters to construct the GEK matrix, becomes advantageous under such conditions.
In contrast, when the angle error exceeds $30^{\circ}$, the QD-SMDS algorithm consistently outperforms SMDS, with the performance gap widening as the angle error increases. These results indicate that as the accuracy of the GEK matrix deteriorates, the role of SVD-based noise suppression through low-rank truncation becomes increasingly important. As such, QD-SMDS —which constructs a rank-$1$ GEK matrix—offers improved robustness to angle errors compared to SMDS, which employs a rank-$3$ GEK matrix.
Fig. 4 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario II. As observed in Fig. 3, SMDS achieves higher localization accuracy when the angle error is up to $\epsilon=20^{\circ}$. However, beyond this point, QD-SMDS significantly outperforms SMDS, and the performance gap becomes more pronounced than in Scenario I. This improvement is attributed to the enhanced robustness of QD-SMDS against measurement errors, enabled by the additional angular information that can be measured or estimated using planar antennas.
IV-C3 Simulation Results With Missing Data
Figure 5: Comparison of localization accuracy between the SMDS and QD-SMDS algorithms with the missing GEK matrix in Scenario II.
We now turn our attention to a more practical scenario in which missing data results in a sparse GEK matrix. In Scenario I, since the SMDS algorithm is initially employed for data supplementation, an effective performance comparison involving matrix completion is not feasible; therefore, this subsection focuses on Scenario II. In our setup, $30\%$ of the entries in both the real-domain GEK matrix $\bm{K}_{\mathrm{r}}$ and the quaternion-domain GEK matrix $\bm{K}_{\mathrm{q}}$ are randomly removed, while preserving matrix symmetry. For SMDS, the incomplete real-domain GEK matrix is first completed using the low-rank matrix completion method described in [31], followed by execution of the SMDS algorithm. For QD-SMDS, the incomplete quaternion-domain GEK matrix is first decomposed into two complex matrices, each of which is individually completed using the same low-rank matrix completion method. These are then recombined into a single quaternion matrix, after which the QD-SMDS algorithm is executed.
Fig. 5 shows the localization accuracy comparison between SMDS and QD-SMDS in Scenario II under missing data conditions. As observed, when the angle error is small (e.g., $\epsilon=10^{\circ}$ and $20^{\circ}$), SMDS still yields the best performance, although the gap between SMDS and QD-SMDS is negligible. In contrast, as the angle error increases, QD-SMDS significantly outperforms SMDS, with the performance gap becoming even more pronounced than in the case without missing data. This improvement is attributed to the lower rank of the quaternion-domain GEK matrix compared to the real-domain one. Since low-rank structures are more favorable for matrix completion, QD-SMDS benefits from improved recovery accuracy, which directly enhances localization performance under larger angle errors.
Based on the numerical results and the computational effort involved, the SMDS algorithm is preferable in scenarios with relatively small angle errors. In contrast, QD-SMDS is more suitable when angle errors are large. Moreover, the supplementary azimuth and elevation angle information obtained from planar antenna arrays becomes increasingly critical as the angle error escalates. Additionally, in the presence of partial angle information loss, QD-SMDS remains effective regardless of the angle error magnitude, demonstrating its robustness to missing data.
V Quaternion-Domain MRC-SMDS
In QD-SMDS, the primary source of computational complexity is the QSVD process. As described in Section II-B, QSVD requires performing SVD on a complex-valued matrix that is twice the size of the original quaternion-domain GEK matrix. To reduce this computational overhead, this section revisits the structural properties of the quaternion-domain GEK matrix and introduces low-complexity variants of the QD-SMDS algorithm that eliminate the need for QSVD.
V-A QD-MRC-SMDS
The quaternion edge vector defined in (28) is partitioned into two sub-vectors: $\bm{\nu}_{\mathrm{AA}}$, corresponding to the edges between ANs, and $\bm{\nu}_{\mathrm{AT}}$, corresponding to the edges between ANs and TNs, as follows:
$\bm{\nu}=\left[\bm{\nu}_{\mathrm{AA}}^{\mathsf{T}},\bm{\nu}_{\mathrm{AT}}^{\mathsf{T}}\right]^{\mathsf{T}},$ (41)
where $\bm{\nu}_{\mathrm{AA}}$ is known since it is fully determined by the coordinates of ANs, and $\bm{\nu}_{\mathrm{AT}}$ can be expressed by
$\bm{\nu}_{\mathrm{AT}}=\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}-\bm{B}_{\mathrm{AT}}\bm{\chi}_{\mathrm{T}},$ (42)
where $\bm{\chi}_{\mathrm{A}}$ and $\bm{\chi}_{\mathrm{T}}$ denotes the quaternion coordinate vectors corresponding to the real-valued coordinate matrices $\bm{X}_{\mathrm{A}}$ and $\bm{X}_{\mathrm{T}}$ for ANs and TNs, respectively, as defined in (14) and (15).
The corresponding structure matrices are given by
$\bm{B}_{\mathrm{AA}}\triangleq\bm{I}_{N_{\mathrm{A}}}\otimes\bm{1}_{N_{\mathrm{T}\times 1}}\quad\text{and}\quad\bm{B}_{\mathrm{AT}}\triangleq\bm{1}_{N_{\mathrm{A}\times 1}}\otimes\bm{I}_{N_{\mathrm{T}}}.$ (43)
From (42), it is evident that the estimation of TN coordinates can be accomplished by accurately estimating the quaternion edge vector $\bm{\nu}_{\mathrm{AT}}$. Furthermore, based on (41), the quaternion-domain GEK matrix in (IV-A1) can be expressed as
$\bm{K}_{\mathrm{q}}=\bm{\nu}\bm{\nu}^{\mathsf{H}}=\!\left[\begin{array}[]{c|c}\bm{\nu}_{\mathrm{AA}}\bm{\nu}_{\mathrm{AA}}^{\mathsf{H}}&\bm{\nu}_{\mathrm{AA}}\bm{\nu}_{\mathrm{AT}}^{\mathsf{H}}\\ \hline\cr\bm{\nu}_{\mathrm{AT}}\bm{\nu}_{\mathrm{AA}}^{\mathsf{H}}&\bm{\nu}_{\mathrm{AT}}\bm{\nu}_{\mathrm{AT}}^{\mathsf{H}}\end{array}\right]\!=\!\left[\begin{array}[]{c|c}\bm{K}_{1}&\bm{K}_{2}\\ \hline\cr\bm{K}_{2}^{\mathsf{H}}&\bm{K}_{3}\end{array}\right].$ (44)
By exploiting the structure observed in (44), the edge vector can be estimated through simple linear filtering without the need for more complex operations.
First, using (42), $\bm{K}_{2}$ can be rewritten as
$\bm{K}_{2}=\bm{\nu}_{\mathrm{AA}}\bm{\nu}^{\mathsf{H}}_{\mathrm{AT}}=\bm{\nu}_{\mathrm{AA}}\left(\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}-\bm{B}_{\mathrm{AT}}\bm{\chi}_{\mathrm{T}}\right)^{\mathsf{H}}.$ (45)
Next, MRC is applied to suppress the influence of $\bm{\nu}_{\mathrm{AA}}$ on the right-hand side, resulting in
$\frac{\bm{\nu}_{\mathrm{AA}}^{\mathsf{H}}}{\|\bm{\nu}_{\mathrm{AA}}\|^{2}}\bm{K}_{2}=\left(\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}-\bm{B}_{\mathrm{AT}}\bm{\chi}_{\mathrm{T}}\right)^{\mathsf{H}}.$ (46)
Finally, taking into account the non-commutativity of both matrix and quaternion multiplications, we solve (46) for $\bm{\chi}_{\mathrm{T}}$, thereby obtaining a closed-form expression for the estimated TN coordinates:
$\bm{\hat{\chi}}_{\mathrm{T}}=\frac{\bm{B}_{\mathrm{AT}}^{\mathsf{T}}}{N_{\mathrm{A}}}\left(\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}-\frac{1}{\|\bm{\nu}_{\mathrm{AA}}\|^{2}}\bm{K}_{2}^{\mathsf{H}}\bm{\nu}_{\mathrm{AA}}\right).$ (47)
This low-complexity variant of the QD-SMDS algorithm is hereafter referred to as QD-MRC-SMDS, with its pseudo-code provided in Algorithm 3. Compared to the original QD-SMDS, QD-MRC-SMDS significantly reduces computational complexity by eliminating expensive operations such as QSVD, matrix inversion, and Procrustes transformation. Instead, it enables direct coordinate estimation via simple quaternion-domain multiplications by utilizing a specific subset of the quaternion-domain GEK matrix.
Algorithm 3 QD-MRC-SMDS
1:Input:
2:Measured and estimated pairwise distances and ADoAs: $\tilde{d}_{m}$, $\tilde{\alpha}_{mp}$, $\tilde{\phi}_{m}^{(\mathrm{xy})}$, $\tilde{\phi}_{m}^{(\mathrm{xz})}$, $\tilde{\phi}_{m}^{(\mathrm{yz})}$, $\tilde{\theta}_{m}^{(\mathrm{x})}$, $\tilde{\theta}_{m}^{(\mathrm{y})}$, $\tilde{\theta}_{m}^{(\mathrm{z})}$
3:The coordinates of all ANs.
4:Steps:
5:Construct $\bm{\nu}_{\mathrm{AA}}$ and $\bm{K}_{2}$ in Eqs. (28) and (IV-A1).
6:Construct $\bm{B}_{\mathrm{AA}}$ and $\bm{B}_{\mathrm{AT}}$ in Eqs. (43).
[…]
8:Convert $\hat{\bm{\chi}}_{\mathrm{T}}$ to the estimated real-valued coordinates matrix $\hat{\bm{X}}_{\mathrm{T}}$.
V-B Iterative QD-MRC-SMDS
While the QD-MRC-SMDS algorithm offers extremely low computational complexity, it leverages only a subset of the information encoded in the GEK matrix and therefore does not fully exploit all available data (i.e., $\bm{K}_{3}$). To overcome this limitation, we extend QD-MRC-SMDS into an iterative estimation algorithm that incorporates the full information contained in the GEK matrix, with only a modest increase in computational cost.
First, from (44), we have
$\begin{bmatrix}\bm{K}_{2}\\ \bm{K}_{3}\end{bmatrix}=\begin{bmatrix}\bm{\nu}_{\mathrm{AA}}\\ \bm{\nu}_{\mathrm{AT}}\end{bmatrix}\bm{\nu}_{\mathrm{AT}}^{\mathsf{H}}.$ (48)
Since $\bm{\nu}_{\mathrm{AT}}$ appears twice in (48), it cannot be uniquely determined in a single step. To address this, we reformulate the problem as an iterative process, from which the following update equation can be derived as
$\bm{\nu}_{\mathrm{AT}}^{(\tau+1)}=\frac{\begin{bmatrix}\bm{K}_{2}^{\mathsf{H}}&\bm{K}_{3}^{\mathsf{H}}\end{bmatrix}}{\left\|\begin{bmatrix}\bm{\nu}_{\mathrm{AA}}^{\mathsf{H}}&(\bm{\nu}_{\mathrm{AT}}^{(\tau)})^{\mathsf{H}}\end{bmatrix}\right\|^{2}}\begin{bmatrix}\bm{\nu}_{\mathrm{AA}}\\ \bm{\nu}_{\mathrm{AT}}^{(\tau)}\end{bmatrix},$ (49)
where $\bm{\nu}_{\mathrm{AT}}^{(\tau)}$ denotes the estimate at the $\tau$-th iteration step ($0\leq\tau\leq\tau_{\mathrm{max}}$).
The initial estimate of $\bm{\nu}_{\mathrm{AT}}$ for the first iteration ($\tau=0$) is given by
$\bm{\nu}_{\mathrm{AT}}^{(0)}=\frac{\bm{K}_{2}^{\mathsf{H}}}{\|\bm{\nu}_{\mathrm{AA}}\|^{2}}\bm{\nu}_{\mathrm{AA}}.$ (50)
Finally, given the final estimate $\bm{\nu}_{\mathrm{AT}}^{(\tau_{\mathrm{max}})}$, the quaternion coordinate vector corresponding to TNs can be obtained from (42) as
$\bm{\hat{\chi}}_{\mathrm{T}}=\frac{\bm{B}_{\mathrm{AT}}^{\mathsf{T}}}{N_{\mathrm{A}}}\left(\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}-\bm{\nu}_{\mathrm{AT}}^{(\tau_{\mathrm{max}})}\right).$ (51)
Algorithm 4 Iterative QD-MRC-SMDS
1:Input:
2:Measured and estimated pairwise distances and ADoAs: $\tilde{d}_{m}$, $\tilde{\alpha}_{mp}$, $\tilde{\phi}_{m}^{(\mathrm{xy})}$, $\tilde{\phi}_{m}^{(\mathrm{xz})}$, $\tilde{\phi}_{m}^{(\mathrm{yz})}$, $\tilde{\theta}_{m}^{(\mathrm{x})}$, $\tilde{\theta}_{m}^{(\mathrm{y})}$, $\tilde{\theta}_{m}^{(\mathrm{z})}$
3:The coordinates of all ANs.
4:Steps:
5:Construct $\bm{\nu}_{\mathrm{AA}}$, $\bm{K}_{2}$ and $\bm{K}_{3}$ via equations in Eqs. (28) and (IV-A1).
6:Initialize $\bm{\nu}_{\mathrm{AT}}$ using Eq. (50).
7:for $\tau=1,\ldots,\tau_{\mathrm{max}}$
8:Update $\bm{\nu}_{\mathrm{AT}}$ using Eq. (49).
9:end for
10:Obtain $\hat{\bm{\chi}}_{T}$ using Eq.(51).
11:Convert $\hat{\bm{\chi}}_{\mathrm{T}}$ to the estimated real-valued coordinates matrix $\hat{\bm{X}}_{\mathrm{T}}$.
This iterative variant of the QD-MRC-SMDS algorithm is hereafter referred to as iterative QD-MRC-SMDS, with its pseudo-code provided in Algorithm 4.
V-C Performance Assessment
V-C1 Iterative Convergence Analysis
Figure 6: Convergence of iterative QD-MRC-SMDS under various ranging and angle measurement errors.
Before comparing the performance of the algorithms, we first examine the convergence behavior of the iterative QD-MRC-SMDS. The matrix $\left[\bm{K}_{2}^{\mathsf{H}}\ \bm{K}_{3}^{\mathsf{H}}\right]$ exhibits a rank-$1$ structure, and (49) is structurally equivalent to the power iteration method. As a result, the direction of $\bm{\nu}_{\mathrm{AT}}$ is expected to converge within a single iteration. Furthermore, since the known vector $\bm{\nu}_{\mathrm{AA}}$ appears explicitly in (49), the scaling of $\bm{\nu}_{\mathrm{AT}}$ is also preserved. Thus, both the direction and scale of $\bm{\nu}_{\mathrm{AT}}$ are anticipated to converge after just one iteration.
Fig. 6 shows the average estimation error of the iterative QD-MRC-SMDS algorithm as a function of the number of iterations $\tau$, under distance error conditions of $\sigma_{\mathrm{d}}=2$ [m] and $4$ [m]. The simulation setup follows the conditions described Section IV-C. As theoretically anticipated, the algorithm converges within a single iteration. This confirms that setting $\tau_{\mathrm{max}}=1$ is sufficient, and the additional computational cost introduced by the iterative procedure corresponds to computing (49) only once.
V-C2 Simulation Results
Under the same simulation conditions as in Figs. 3 and 4, the localization accuracy of the QD-MRC-SMDS and iterative QD-MRC-SMDS algorithms is compared.
Fig. 7 shows the localization performance of each method in Scenario I, with the QD-SMDS algorithm included as reference (black curve). Notably, both QD-MRC-SMDS and its iterative variant achieve high localization accuracy without relying on SVD-based noise suppression. Even the lowest-complexity QD-MRC-SMDS incurs only a modest performance degradation compared to QD-SMDS. Moreover, its ability to maintain high accuracy under large angle errors indicates that QD-MRC-SMDS inherits the robustness to angular uncertainties observed in QD-SMDS. When compared to the SMDS performance shown in Fig. 3, QD-MRC-SMDS demonstrates superior localization accuracy, especially severe angular error conditions (i.e., $\epsilon\geq 30^{\circ}$), underscoring its robustness and computational efficiency. Furthermore, the iterative QD-MRC-SMDS algorithm achieves performance nearly equivalent to that of QD-SMDS, suggesting that the full potential of the GEK matrix can be exploited without requiring low-rank truncation via SVD.
Figure 7: Comparison of localization accuracy between the QD-SMDS, QD-MRC-SMDS, and iterative QD-MRC-SMDS algorithms in Scenario I.
Fig. 8 shows the localization performance of each method in Scenario II. While the overall trend is similar to that observed in Scenario I, the performance gap between the iterative QD-MRC-SMDS and QD-SMDS increases with larger angular errors. This can be attributed primarily to the greater degradation in the accuracy of the GEK matrix caused by increased measurement errors in the additional angular parameters of the azimuth and elevation angles. In this case, it appears that noise suppression in the SVD-based approach works slightly better than in the iterative one. Nevertheless, the performance gap remains minimal—even at the largest angular error ($\epsilon=50^{\circ}$), the estimation error difference is within $0.05$ [m].
Figure 8: Comparison of localization accuracy between the QD-SMDS, QD-MRC-SMDS, and iterative QD-MRC-SMDS algorithms in Scenario II.
Based on these results, system designers can select the most appropriate algorithm by balancing the computational complexity against localization accuracy. For latency-critical applications, the low-complexity QD-MRC-SMDS offers the most practical choice. Conversely, when computational resources are sufficient and maximum accuracy is required—even under large angular errors—QD-SMDS becomes the preferred option. Lastly, the iterative QD-MRC-SMDS achieves an excellent trade-off between accuracy and complexity, making it a strong candidate for a wide range of practical deployment scenarios.
VI Conclusion
In this paper, we proposed a novel QD-SMDS algorithm, developed by reformulating the classical SMDS algorithm within the quaternion domain to enable low-complexity, simultaneous localization of multiple targets using data aggregated from a large number of wireless sensor nodes. By constructing the GEK matrix in the quaternion domain, the matrix rank can be reduced to one, even in 3D Euclidean space, thereby maximizing the noise suppression effect via QSVD. Moreover, the proposed method is inherently compatible with low-rank matrix completion techniques, which further enhances its robustness against missing data. Simulation results demonstrate that QD-SMDS consistently outperforms the conventional SMDS, particularly in scenarios with significant angle measurement errors. Its advantage becomes more pronounced when both azimuth and elevation information are available. However, the use of QSVD —along with matrix inversion and the Procrustes transformation required for coordinate recovery—introduces substantial computational complexity. To address this limitation, we further developed a computationally efficient variants of QD-SMDS by exploiting the structural properties of the quaternion-domain GEK matrix. These variants enable coordinate estimation through simple quaternion matrix multiplications, achieving localization performance comparable to that of the original QD-SMDS, while significantly reducing computational cost.